# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v1

@@ 1 Introduction
Figure 2: An overview diagram illustrating the structure of this survey.
@@ 2.1.1 Application Tasks Category
$\bullet$ 3DGS Segmentation. Based on the input space $\mathcal{X}$, existing 3DGS segmentation methods can be broadly classified into two categories: generic segmentation and promptable segmentation. Generic segmentation aims to predict pixel- or point-level labels without external prompts, and can be further divided into: (i) Semantic segmentation [17]; (ii) Instance segmentation [18, 19, 20]; and (iii) Panoptic segmentation [21, 22]. Promptable segmentation extends the generic setting by introducing external prompts to guide the segmentation process. Depending on the form of prompts, this category includes: (i) Interactive segmentation, where users provide spatial hints such as clicks, scribbles, boxes, or masks [23, 24, 25]; (ii) Open-vocabulary segmentation, which uses category names as textual prompts to support segmenting novel or unseen classes [26, 27]; (iii) Referring segmentation, which relies on natural language expressions to localize and segment specific objects [7]. In addition, depending on the representation of the output space $\mathcal{Y}$, segmentation results can be produced either on rendered 2D images [26, 28] or directly over the 3D Gaussian representations [29, 30, 31, 32].
@@ 2.1.2 Learning Paradigms for Application Tasks
where $\varepsilon(\cdot,\cdot)$ denotes a task-specific reconstruction loss function, such as an $\ell_{1}$ loss, depending on the target application. This per-scene fitting scheme enables high-fidelity modeling tailored to each scene, making it well-suited for tasks requiring view-consistent reconstruction or dense appearance modeling. However, the lack of generalization and high computational cost make it impractical for large-scale or real-time applications.
@@
$\bullet$ Diffusion Models (DMs). These generative models start from Gaussian noise and progressively denoise it, step-by-step, into a coherent image; the denoiser is trained to invert a fixed noise schedule (DDPM [54]). Latent diffusion (LDM [55]) performs the same procedure in the auto-encoder’s latent space, making high-resolution synthesis tractable on everyday GPUs. Text-conditioned variants, e.g., Stable Diffusion [55], inject cross-attention to a language encoder, allowing precise, prompt-driven control while preserving visual fidelity. Their attention maps have been shown to correlate with meaningful regions in the output, connecting image generation with vision–language understanding.
@@ 2.3 Related Research Areas
$\bullet$ Segmentation Based on NeRF. Semantic NeRF [56] pioneers the incorporation of semantic information into radiance fields by introducing per-point semantic layers. Follow-up works can be broadly categorized into two directions. The first focuses on feature distillation, where high-dimensional 2D features from pretrained vision models (e.g., DINO, CLIP) are embedded into 3D neural fields. Representative methods such as Distilled Feature Fields [57], Neural Feature Fusion Fields [58], ISRF [59], LERF [60], and 3D-OVS [61] optimize volumetric feature fields to reconstruct semantic features via differentiable rendering. The second line of research targets mask lifting, which seeks to overcome the scarcity of 3D annotations by projecting 2D semantic masks into 3D space. Early methods rely on either ground-truth masks [62, 63] or pretrained segmentation models, such as methods like Panoptic Lifting [64], ContrastiveLift [65], and SPIn-NeRF [66]. However, these approaches often suffer from limited generalization due to closed-set assumptions. With the recent emergence of powerful foundation models like SAM and SAM2 [48], newer works aim to leverage their open-world segmentation capabilities. For instance, SA3D [67] proposes an interactive pipeline that propagates a single SAM mask across multiple views into a NeRF-based scene, enabling sparse-to-dense segmentation. GARField [68] further explores integrating SAM-derived cues into radiance field optimization.
@@ 3.1.1 Feature Distillation-based Methods
These methods [26, 27, 83, 84] aim to distill the semantic knowledge embedded in 2D foundation models (e.g., CLIP[47], SAM [48], DINO [45]) into 3D scene representations. By leveraging the semantic understanding learned from large-scale 2D datasets, these approaches enhance 3D models with open-vocabulary recognition capabilities. LangSplat [26] and Feature3DGS [27] are among the earliest works to explore this direction. LangSplat first segments images using SAM [48], and then feeds the resulting hierarchical semantic masks into CLIP to extract region-level semantic embeddings. To alleviate memory constraints, it employs a scene-specific language autoencoder to compress these high-dimensional semantic features. These compressed features are used to supervise a set of initialized 3D Gaussian language features, effectively transferring CLIP’s open-vocabulary semantics into the 3D domain for downstream tasks. Feature3DGS, developed concurrently, distills features from LSeg [85] and SAM [48], and further employs SAM’s decoder to interpret 2D rendered views, establishing a bridge between 2D and 3D segmentation. While effective, these methods may suffer from semantic information loss and high computational cost. To address these limitations, subsequent works have evolved along two key directions: improving segmentation accuracy and reducing computation overhead.
@@
$\bullet$ Enhancing Segmentation Accuracy. N2F2 [86] addresses LangSplat’s inefficiency in scale selection by learning a unified high-dimensional feature field under hierarchical supervision, where each feature dimension encodes scene semantics at different granularities, enabling fine-grained yet efficient representation without multi-scale querying. LangSurf [87] improves LangSplat by extracting dense pixel-level semantics from the entire image, followed by SAM-based mask pooling to preserve contextual cues. It also jointly optimizes language Gaussians on object surfaces using geometry supervision and contrastive loss, enhancing object’s spatial consistency in 3D. seconGS [88] further refines LangSplat by using SAM2-generated masklets to provide semantically consistent supervision. To overcome the limitations of 2D querying in 3D space, it introduces a two-step querying mechanism that first retrieves distilled ground-truth and then queries individual Gaussians for spatially accurate 3D understanding. VLGaussian [28] observes that prior color-based rasterization is unsuitable for language modalities due to its dependence on the same color opacity. Instead, it introduces a cross-modal rasterizer tailored for rendering semantic language features with newly introduced opacity for language. SuperGSeg [89], inspired by superpoint representations, clusters Gaussians into Super-Gaussians using spatial and instance cues based on Scaffold-GS [3]. High-dimensional language features are assigned to each Super-Gaussian to support structured and comprehensive scene understanding while preserving linguistic granularity. Unlike existing methods that distill 2D knowledge to 3D projected view, SemanticGaussians [90] design a versatile projection approach that maps various 2D semantic features from pre-trained image encoders into a semantic component of 3D Gaussians based on spatial relationships. It distills projected 2D knowledge into 3D semantic network directly for fast inference. Beyond the aforementioned approaches that solely introduce semantic information, GLS [91] incorporates geometric priors (surface normals) to effectively refine the rendered depth.
@@
$\bullet$ Consistency-Oriented Pre-Processing. Several methods [22, 19, 21, 98, 99, 100] aim to improve multi-view mask consistency through pre-processing strategies before lifting 2D masks into 3D. GaussianGrouping [22] introduces an object association technique as a pre-processing step to align 2D segmentation maps across views, enhancing cross-view consistency. However, the quality of such pre-processing is often limited by errors in object correspondence under varying viewpoints. Gaga [19] addresses this by leveraging spatial cues and a 3D-aware memory bank to effectively associate object masks across diverse camera poses. CCGS [21] further argues that Gaga lacks segmentation-aware optimization within the Gaussian representation itself, which may lead to inconsistent or floating structures in the 3D segmentation field. To overcome this, CCGS constructs a unified 3D point cloud field and performs segmentation-constrained optimization to reinforce mask association through spatial regularization.
@@
$\bullet$ Consistency-Aware End-to-End Learning. To reduce reliance on heuristic pre- and post-processing, recent works [30, 102, 24, 32, 103, 17, 104] propose end-to-end learning frameworks that directly construct consistent and distinguishable 3D feature fields. SAGA [30] distills 2D masks into 3D features that introduce a scale-aware contrastive training strategy that distills SAM’s segmentation capability into a scale-gated affinity representation, which handles multi-granularity ambiguity. SA3D [102] introduces an iterative refinement strategy that alternates between inverse rendering of 2D masks and cross-view self-prompting, enabling progressively more consistent and accurate object segmentation across multiple views. Click-Gaussian [24] enables efficient interactive segmentation of 3D Gaussians by incorporating multi-granularity feature fields derived from 2D masks. It further introduces Global Feature-guided Learning (GFL) to alleviate inconsistencies across different views. Unified-Lift [20] develops an object-aware lifting pipeline that eliminates the need for additional alignment steps. It formulates a codebook-based object representation and aligns object-level semantics with Gaussian features via contrastive learning, leading to improved 3D instance understanding. OpenGaussian [32] focuses on 3D open-vocabulary segmentation by leveraging SAM-predicted masks to supervise 3D instance features. It adopts a coarse-to-fine feature discretization strategy via a two-stage codebook and introduces an instance-level 3D–2D association module that bridges Gaussian points with 2D masks and CLIP embeddings. econSG [105] improves zero-shot 3D semantic understanding by projecting refined multi-view language features into a low-dimensional 3D latent space, enabling efficient optimization and consistent semantic field initialization within the 3D Gaussian framework. CCL-LGS [106] aligns SAM-generated masks across views using a zero-shot tracker and extracts semantic features with CLIP. A contrastive codebook learning module is then applied to enforce cross-view consistency by promoting intra-class compactness and inter-class separation in the learned semantic space. VoteSplat [107] integrates Hough voting with 3DGS by learning a per-Gaussian 3D offset whose projections align with SAM-derived 2D vote maps and are depth-regularized, so that votes aggregate near object centroids. ILGS [108] addresses view-inconsistent language embeddings challenge by introducing two key components: an identity-aware semantic consistency loss that enforces alignment of language features across views, and a progressive mask expansion to enhance boundary precision by modeling local segment relationships.
@@ 3.1.3 Feed-Forward-based Methods
To overcome the inefficiency of per-scene optimization, especially under dense calibrated images conditions, several methods [115, 116, 117, 118, 119] adopt feed-forward architectures to enable fast and generalizable 3D semantic field construction. SLGaussian [115] builds a feed-forward pipeline that infers 3D semantic Gaussians directly from sparse viewpoints based on MVSplat [4]. It enforces mask consistency via video-based tracking and embeds language features by indexing high-dimensional CLIP representations through compact low-dimensional codes, achieving efficient and scalable 3D understanding. Dr. Splat [117] departs from traditional rendering-based supervision by directly associating CLIP embeddings with 3D Gaussians. It introduces a language feature registration mechanism that assigns CLIP features to dominant Gaussians intersected by pixel rays. Additionally, it employs Product Quantization (PQ) learned from large-scale image datasets to compress features. LangScene-X [119] first uses a TriMap video diffusion network based on CogVideoX [120] to densify sparse inputs, synthesizing appearance, geometry, and semantics. It then employs a Language-Quantized Compressor (LQC) to encode language embeddings on large-scale datasets. SceneSplat [121] introduces a self-supervised framework that facilitates rich 3D feature learning from unlabeled scenes. To support this effort, SceneSplat-7K is introduced as the first large-scale dataset specifically designed for 3DGS in indoor environments.
@@ 3.2.1 Editing Based on Text Prompt
$\bullet$ Efforts towards Multi-View Consistency. Editing 3D objects while maintaining multi-view consistency is a core challenge in the field of 3D editing. GaussCtrl [130] leverages depth-conditioned editing by utilizing naturally consistent depth maps to ensure geometric consistency across multi-view images, along with attention-based latent code alignment to achieve a unified appearance across views. Luo et al. [131] propose a progressive 3D editing strategy using a Trajectory-Anchored Scheme (TAS) and a dual-branch editing mechanism to ensure multi-view consistency, addressing challenges in error accumulation during text-to-image processes. Eyal et al. [132] introduce a geometry-guided warping mechanism that leverages the scene’s depth and structural information to accurately map edits across views, ensuring multi-view consistency. DGE [133] employs spatiotemporal attention to jointly edit the selected key views and inject their features into other views, using correspondences derived from epipolar-constrained visual features, thereby establishing multi-view consistency. SplatFlow [134] leverages a multi-view rectified flow (RF) model to simultaneously generate multi-view images, depth information, and camera poses based on text prompts. By incorporating training-free inversion and inpainting techniques, it achieves seamless editing of 3DGS. InterGSEdit [135] introduces a 3D Geometry-Consistent Attention Prior and an Adaptive Cross-Dimensional Attention Fusion Network to ensure multi-view consistency and enable fine-grained detail recovery.
$\bullet$ Efforts towards Efficiency and Speed. Additionally, to enhance the efficiency and speed of 3D editing, ProGDF [136] comprises Progressive Gaussian Sampling (PGS) and Gaussian Difference Field (GDF). PGS applies progressive constraints to generate diverse intermediate results during the editing process. GDF employs a lightweight neural network, leveraging these intermediate results to model the editing process, allowing for real-time, controllable, and flexible editing within a single training session. The optimization process of DreamCatalyst [36] approximates the reverse diffusion process, aligning with diffusion sampling dynamics and thereby reducing training time. 3DSceneEditor [137] utilizes a streamlined 3D pipeline, allowing direct manipulation of Gaussians. By leveraging input prompts, semantic labeling, and CLIP’s zero-shot capabilities, it enables efficient, high-quality editing. Chen et al. [133] propose Direct Gaussian Editor, which adapts 2D image editing models into a multi-view consistent version and integrates the 3D geometric representation of the underlying scene. This method enables direct optimization of the 3D representation, eliminating the need for iterative editing. 3DitScene [138] incorporates generative priors and optimization techniques to refine the 3D Gaussian model and utilizes language features extracted by CLIP for object disentanglement.

### v2

@@ 1 Introduction
Figure 2: Overview of the structure of this survey.
@@ 2.1.1 Application Tasks Category
$\bullet$ 3DGS Segmentation. Based on the input space $\mathcal{X}$, existing 3DGS segmentation methods can be broadly classified into two categories: generic segmentation and promptable segmentation. Generic segmentation aims to predict pixel- or point-level labels without external input prompts, and can be further divided into: (i) Semantic segmentation [17]; (ii) Instance segmentation [18, 19, 20]; and (iii) Panoptic segmentation [21, 22]. Promptable segmentation extends the generic setting by introducing external input prompts to guide the segmentation process. Depending on the form of prompts, this category includes: (i) Interactive segmentation, where users provide spatial hints such as clicks, scribbles, boxes, or masks [23, 24, 25]; (ii) Open-vocabulary segmentation, which uses category names as textual prompts to support segmenting novel or unseen classes [26, 27]; (iii) Referring segmentation, which relies on natural language expressions to localize and segment specific objects [7]. In addition, depending on the representation of the output space $\mathcal{Y}$, segmentation results can be produced either on rendered 2D images [26, 28] or directly over the 3D Gaussian representations [29, 30, 31, 32].
@@ 2.1.2 Learning Paradigms for Application Tasks
where $\varepsilon(\cdot,\cdot)$ denotes a task-specific reconstruction loss function, such as an $\ell_{1}$ loss, depending on the target application. This per-scene learning scheme enables high-fidelity modeling tailored to each scene, making it well-suited for tasks requiring view-consistent reconstruction or dense appearance modeling. However, the lack of generalization and high computational cost make it impractical for large-scale or real-time applications.
@@
$\bullet$ Diffusion Models (DMs). These generative models start from Gaussian noise and progressively denoise it into an image (DDPM [54]). Latent Diffusion Models [55] (LDM) improve computational efficiency by performing denoising in a compressed latent space, while Stable Diffusion [55] builds on this approach by enabling text-conditioned generation through a CLIP [47] text encoder and a cross-attention mechanism. DiT [56] replaces the U-Net backbone with a Transformer, enhancing scalability and performance, and Flux [57] further advances this line by adopting a flow-matching framework with Transformer backbones, enabling faster sampling and higher-fidelity image generation.
@@ 2.3 Related Research Areas
$\bullet$ Segmentation Based on NeRF. Semantic NeRF [58] pioneers the incorporation of semantic information into radiance fields by introducing per-point semantic layers. Follow-up works can be broadly categorized into two directions. The first focuses on feature distillation, where high-dimensional 2D features from pretrained vision models (e.g., DINO, CLIP) are embedded into 3D neural fields. Representative methods such as Distilled Feature Fields [59], Neural Feature Fusion Fields [60], ISRF [61], LERF [62], and 3D-OVS [63] optimize volumetric feature fields to reconstruct semantic features via differentiable rendering. The second line of research targets mask lifting, which seeks to overcome the scarcity of 3D annotations by projecting 2D semantic masks into 3D space. Early methods rely on either ground-truth masks [64] or pretrained segmentation models, such as methods like Panoptic Lifting [65], ContrastiveLift [66], and SPIn-NeRF [67]. However, these approaches often suffer from limited generalization due to closed-set assumptions. With the recent emergence of powerful foundation models like SAM and SAM2 [48], newer works aim to leverage their open-world segmentation capabilities. For instance, SA3D [68] proposes an interactive pipeline that propagates a single SAM mask across multiple views into a NeRF-based scene, enabling sparse-to-dense segmentation. GARField [69] integrates SAM-derived masks into radiance field by learning a scale-conditioned affinity field.
@@ 3.1.1 Feature Distillation-based Methods
These methods [26, 27, 84, 85, 86] aim to distill the semantic knowledge embedded in 2D foundation models (e.g., CLIP[47], SAM [48], DINO [45]) into 3D scene representations. By leveraging the semantic understanding learned from large-scale 2D datasets, these approaches enhance 3D models with open-vocabulary recognition capabilities. LangSplat [26] and Feature3DGS [27] are among the earliest works to explore this direction. LangSplat first segments images using SAM [48], and then feeds the resulting hierarchical semantic masks into CLIP to extract region-level semantic embeddings. To alleviate memory constraints, it employs a scene-specific language autoencoder to compress these high-dimensional semantic features. These compressed features are used to supervise a set of initialized 3D Gaussian language features, effectively transferring CLIP’s open-vocabulary semantics into the 3D domain for downstream tasks. Feature3DGS, developed concurrently, distills features from LSeg [87] and SAM [48], and further employs SAM’s decoder to interpret 2D rendered views, establishing a bridge between 2D and 3D segmentation. While effective, these methods may suffer from semantic information loss and high computational cost. To address these limitations, subsequent works have evolved along two key directions: improving segmentation accuracy and reducing computation overhead.
@@
$\bullet$ Enhancing Segmentation Accuracy. N2F2 [88] addresses LangSplat’s inefficiency in scale selection by learning a unified high-dimensional feature field under hierarchical supervision, where each feature dimension encodes scene semantics at different granularities, enabling fine-grained yet efficient representation without multi-scale querying. LangSurf [89] improves LangSplat by extracting dense pixel-level semantics from the entire image, followed by SAM-based mask pooling to preserve contextual cues. It also jointly optimizes language Gaussians on object surfaces using geometry supervision and contrastive loss, enhancing object’s spatial consistency in 3D. seconGS [90] further refines LangSplat by using SAM2-generated masklets to provide semantically consistent supervision. To overcome the limitations of 2D querying in 3D space, it introduces a two-step querying mechanism that first retrieves distilled ground-truth and then queries individual Gaussians for spatially accurate 3D understanding. VLGaussian [28] observes that prior color-based rasterization is unsuitable for language modalities due to its dependence on the same color opacity. Instead, it introduces a cross-modal rasterizer tailored for rendering semantic language features with newly introduced opacity for language. SuperGSeg [91], inspired by superpoint representations, clusters Gaussians into Super-Gaussians using spatial and instance cues based on Scaffold-GS [3]. High-dimensional language features are assigned to each Super-Gaussian to support structured and comprehensive scene understanding while preserving linguistic granularity. FreeGS [92] introduces a semantic-embedded 3DGS framework by coupling semantic features with instance identities via the proposed IDSF field. A two-step alternating optimization and a 2D–3D contrastive loss ensure view-consistent semantics without requiring 2D labels. Beyond the aforementioned approaches that solely introduce semantic information, GLS [93] incorporates geometric priors (surface normals) to effectively refine the rendered depth.
@@
$\bullet$ Consistency-Oriented Pre-Processing. Several methods [22, 19, 21, 100, 101, 102] aim to improve multi-view mask consistency through pre-processing strategies before lifting 2D masks into 3D. GaussianGrouping [22] introduces an object association technique as a pre-processing step to align 2D segmentation maps across views, enhancing cross-view consistency. However, the quality of such pre-processing is often limited by errors in object correspondence under varying viewpoints. Gaga [19] addresses this by leveraging spatial cues and a 3D-aware memory bank to effectively associate object masks across diverse camera poses. CCGS [21] further argues that Gaga lacks segmentation-aware optimization within the Gaussian representation itself, which may lead to inconsistent or floating structures in the 3D segmentation field. To overcome this, CCGS constructs a unified 3D point cloud field and performs segmentation-constrained optimization to reinforce mask association.
@@
$\bullet$ Consistency-Aware End-to-End Learning. To reduce reliance on heuristic pre- and post-processing, recent works [30, 104, 24, 32, 105, 17, 106] propose end-to-end learning frameworks that directly construct consistent and distinguishable 3D feature fields. SAGA [30] distills 2D masks into 3D features that introduce a scale-aware contrastive training strategy that distills SAM’s segmentation capability into a scale-gated affinity representation, which handles multi-granularity ambiguity. SA3D [104] introduces an iterative refinement strategy that alternates between inverse rendering of 2D masks and cross-view self-prompting, enabling progressively more consistent and accurate object segmentation across multiple views. Click-Gaussian [24] enables efficient interactive segmentation of 3D Gaussians by incorporating multi-granularity feature fields derived from 2D masks. It further introduces Global Feature-guided Learning (GFL) to alleviate inconsistencies across different views. Unified-Lift [20] develops an object-aware lifting pipeline that eliminates the need for additional alignment steps. It formulates a codebook-based object representation and aligns object-level semantics with Gaussian features via contrastive learning, leading to improved 3D instance understanding. OpenGaussian [32] focuses on 3D open-vocabulary segmentation by leveraging SAM-predicted masks to supervise 3D instance features. It adopts a coarse-to-fine feature discretization strategy via a two-stage codebook and introduces an instance-level 3D–2D association module that bridges Gaussian points with 2D masks and CLIP embeddings. econSG [107] improves zero-shot 3D semantic understanding by projecting refined multi-view language features into a low-dimensional 3D latent space, enabling efficient optimization and consistent semantic field initialization within the 3D Gaussian framework. CCL-LGS [108] aligns SAM-generated masks across views via a zero-shot tracker, extracts semantics with CLIP, and applies contrastive codebook learning to enforce cross-view consistency by promoting intra-class compactness and inter-class separation. VoteSplat [109] integrates Hough voting with 3DGS by learning per-Gaussian 3D offsets whose projections align with SAM-derived 2D vote maps and are depth-regularized, so that votes aggregate near object centroids. ILGS [110] mitigates view-inconsistent language embeddings by enforcing cross-view semantic alignment with an identity-aware semantic consistency loss and refining boundaries via progressive mask expansion.
@@ 3.1.3 Feed-Forward-based Methods
To overcome the inefficiency of per-scene optimization, especially under dense calibrated images conditions, several methods [117, 118, 119, 120, 121, 122] adopt feed-forward architectures to enable fast and generalizable 3D semantic field construction. LSM [123] is the first feed-forward framework that integrates DUSt3R [124], point transformer [125], and LSeg [87] for language-driven scene understanding. SLGaussian [117] builds a feed-forward pipeline that infers 3D semantic Gaussians directly from sparse viewpoints based on MVSplat [4]. Dr. Splat [119] departs from traditional rendering-based supervision by directly associating CLIP embeddings with 3D Gaussians. It introduces a language feature registration mechanism that assigns CLIP features to dominant Gaussians intersected by pixel rays. Additionally, it employs Product Quantization (PQ) learned from large-scale image datasets to compress features. LangScene-X [121] first uses a TriMap video diffusion network based on CogVideoX [126] to densify sparse inputs, synthesizing appearance, geometry, and semantics. It then employs a Language-Quantized Compressor (LQC) to encode language embeddings on large-scale datasets. SceneSplat [127] introduces a self-supervised framework that facilitates rich 3D feature learning from unlabeled scenes. To support this effort, SceneSplat-7K is introduced as the first large-scale dataset specifically designed for 3DGS in indoor environments.
@@ 3.2.1 Editing Based on Text Prompt
$\bullet$ Efforts toward Multi-View Consistency. Editing 3D objects while maintaining multi-view consistency is a core challenge in the field of 3D editing. GaussCtrl [136] leverages depth-conditioned editing by utilizing naturally consistent depth maps to ensure geometric consistency across multi-view images, along with attention-based latent code alignment to achieve a unified appearance across views. Luo et al. [137] propose a progressive 3D editing strategy using a Trajectory-Anchored Scheme (TAS) and a dual-branch editing mechanism to ensure multi-view consistency, addressing challenges in error accumulation during text-to-image processes. Gomel et al. [138] introduce a geometry-guided warping mechanism that leverages the scene’s depth and structural information to accurately map edits across views, ensuring multi-view consistency. DGE [139] employs spatiotemporal attention to jointly edit the selected key views and inject their features into other views, using correspondences derived from epipolar-constrained visual features, thereby establishing multi-view consistency. SplatFlow [140] leverages a multi-view rectified flow (RF) model to simultaneously generate multi-view images, depth information, and camera poses based on text prompts. By incorporating training-free inversion and inpainting techniques, it achieves seamless editing of 3DGS. InterGSEdit [141] introduces a 3D Geometry-Consistent Attention Prior and an Adaptive Cross-Dimensional Attention Fusion Network to ensure multi-view consistency and enable fine-grained detail recovery.
$\bullet$ Efforts toward Efficiency and Speed. Additionally, to enhance the efficiency and speed of 3D editing, ProGDF [142] comprises Progressive Gaussian Sampling (PGS) and Gaussian Difference Field (GDF). PGS applies progressive constraints to generate diverse intermediate results during the editing process. GDF employs a lightweight neural network, leveraging these intermediate results to model the editing process, allowing for real-time, controllable, and flexible editing within a single training session. The optimization process of DreamCatalyst [36] approximates the reverse diffusion process, aligning with diffusion sampling dynamics and thereby reducing training time. 3DSceneEditor [143] utilizes a streamlined 3D pipeline, allowing direct manipulation of Gaussians. By leveraging input prompts, semantic labeling, and CLIP’s zero-shot capabilities, it enables efficient, high-quality editing. Chen et al. [139] propose Direct Gaussian Editor, which adapts 2D image editing models into a multi-view consistent version and integrates the 3D geometric representation of the underlying scene. This method enables direct optimization of the 3D representation, eliminating the need for iterative editing. 3DitScene [144] incorporates generative priors and optimization techniques to refine the 3D Gaussian model and utilizes language features extracted by CLIP for object disentanglement.